40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m elog  Not logged in ELOG logo
Entry  Tue Apr 23 18:48:25 2019, Jon, Update, SUS, c1susaux bench testing completed 
    Reply  Wed Apr 24 17:07:39 2019, gautam, Update, SUS, c1susaux in-situ testing [and future of IFOtest] 2019-04-24_20-29.pdfScreenshot_from_2019-04-24_20-05-54.pngSRM_OSEMPD_WHT_ACROMAG.pdfDCVmon.png
Message ID: 14567     Entry time: Wed Apr 24 17:07:39 2019     In reply to: 14563
Author: gautam 
Type: Update 
Category: SUS 
Subject: c1susaux in-situ testing [and future of IFOtest] 

[jon, gautam]

For the in-situ test, I decided that we will use the physical SRM to test the c1susaux Acromag replacement crate functionality for all 8 optics (PRM, BS, ITMX, ITMY, SRM, MC1, MC2, MC3). To facilitate this, I moved the backplane connector of the SRM SUS PD whitening board from the P1 connector to P2, per Koji's mods at ~5:10PM local time. Watchdog was shutdown, and the backplane connectors for the SRM coil driver board was also disconnected (this is interfaced now to the Acromag chassis).

I had to remove the backplane connector for the BS coil driver board in order to have access to the SRM backplane connector. Room in the back of these eurocrate boxes is tight in the existing config...

At ~6pm, I manually powered down c1susaux (as I did not know of any way to turn off the EPICS server run by the old VME crate in a software way). The point was to be able to easily interface with the MEDM screens. So the slow channels prefixed C1:SUS-* are now being served by the Supermicro called c1susaux2.

A critical wiring error was found. The channel mapping prepared by Johannes lists the watchdog enable BIO channels as "C1:SUS-<OPTIC>_<COIL>_ENABLE", which go to pins 23A-27A on the P1 connector, with returns on the corresponding C pins. However, we use the "TEST" inputs of the coil driver boards for sending in the FAST actuation signals. The correct BIO channels for switching this input is actually "C1:SUS-<OPTIC>_<COIL>_TEST", which go to pins 28A-32A on the P1 connector. For todays tests, I voted to fix this inside the Acromag crate for the SRM channels, and do our tests. Chub will unfortunately have to fix the remaining 7 optics, see Attachment #1 for the corrections required. I apportion 70% of the blame to Johannes for the wrong channel assignment, and accept 30% for not checking it myself.

The good news: the tests for the SRM channels all passed!

  • Attachment #2: Output of Jon's testing code. My contribution is the colored logs courtesy of python's coloredlogs package, but this needs a bit more work - mainly the PASS mssage needs to be green. This test applies bias voltages to PIT/YAW, and looks for the response in the PDmon channels. It backs out the correct signs for the four PDs based on the PIT/YAW actuation matrix, and checks that the optic has moved "sufficiently" for the applied bias. You can also see that the PD signals move with consistent signs when PIT/YAW misalignment is applied. Additionally, the DC values of the PDMon channels reported by the Acromag system are close to what they were using the VME system. I propose calling the next iteration of IFOtest "Sherlock".
  • Attachment #3: Confirmation (via spectra) that the SRM OSEM PD whitening can still be switched even after my move of the signals from the P1 connector to the P2 connector. I don't have an explanation right now for the shape of the SIDE coil spectrum.
  • Attachment #4: Applied 100 cts (~ 100*10/2**15/2 ~ 15mV at the monitor point) offset at the bias input of the coil output filters on SRM (this is a fast channel). Looked for the response in the Coil Vmon channels (these are SLOW channels). The correct coil showed consistent response across all 5 channels.

Additionally, I confirmed that the watchdog tripped when the RMS OSEM PD voltage exceeded 200 counts. Ideally we'd have liked to test the stability of the EPICS server, but we have shut it down and brought the crate back out to the electronics bench for Chub to work on tomorrow.

I restarted the old VME c1susaux at 915pm local time as I didn't want to leave the watchdogs in an undefined state. Unsurprisingly, ITMY is stuck. Also, the BS (cable #22) and SRM (cable #40) coil drivers are physically disconnected at the front DB15 output because of the undefined backplane inputs. I also re-opened the PSL shutter.

Attachment 1: 2019-04-24_20-29.pdf  654 kB  Uploaded Wed Apr 24 21:42:12 2019  | Hide | Hide all | Show all
2019-04-24_20-29.pdf
Attachment 2: Screenshot_from_2019-04-24_20-05-54.png  1.814 MB  Uploaded Wed Apr 24 21:53:49 2019  | Hide | Hide all | Show all
Screenshot_from_2019-04-24_20-05-54.png
Attachment 3: SRM_OSEMPD_WHT_ACROMAG.pdf  53 kB  Uploaded Wed Apr 24 21:54:03 2019  | Hide | Hide all | Show all
SRM_OSEMPD_WHT_ACROMAG.pdf
Attachment 4: DCVmon.png  96 kB  Uploaded Wed Apr 24 21:54:19 2019  | Show | Hide all | Show all
ELOG V3.1.3-